OpenStudy (anonymous):

5(b+2)+5b=-20 then b=____

OpenStudy (anonymous):

go to math section

OpenStudy (anonymous):

4, it's definitely 4

OpenStudy (anonymous):

b=3

OpenStudy (anonymous):

3

OpenStudy (anonymous):

5b + 10 + 5b = -20 10b = -30 b = -3

OpenStudy (anonymous):

believe me - i used to teach maths

OpenStudy (anonymous):

I pretty sure Jimmyrep answer is right. I came to the same result using same steps

OpenStudy (anonymous):

Jimmy rep is correct. I just worked it out and got -3. Here's my steps along with an explanation: 5(b+2)+5b=-20 (original equation) 5b + 10 + 5b = -20 (distributed the 5 to the b and to the 2) 10b + 10 = -20 (combined the b's) 10b = -30 (subtracted 10 from both sides) b=-3 (divided both sides by 10)